OpenStudy (anonymous):

How to factor 2x^2+3x

OpenStudy (anonymous):

You can factor by x. Which you would get: x(2x + 3) Which you would distribute the x to all the other numbers. Roots: - 3/2 and 0.

OpenStudy (anonymous):

There are other forms of that equation.

OpenStudy (anonymous):

thats perfect thanks!

OpenStudy (anonymous):

Welcome. If you need any information on factoring or another question. I may be able to help.

OpenStudy (anonymous):

actually, im having trouble with this one also; 2x^2-13x+20=0

OpenStudy (anonymous):

If you wouldnt mind helping me that would be great

OpenStudy (anonymous):

I think I just did that on my exam, brb let me check.

OpenStudy (anonymous):

Im writing mine tomorrow lol

OpenStudy (anonymous):

(x - 4)(2x - 5) = 0 Where the solutions are: 5/2 and 4. You would plug in those to get 0. and foil those two equations together to get 2x^2-13x+20=0

OpenStudy (anonymous):

how did you get (x - 4)(2x - 5) = 0?

OpenStudy (anonymous):

You break the equation apart. Factor it yourself and you should get 2x^2-13x+20=0

OpenStudy (anonymous):

ok i see what u mean but how did you get from 2x^2-13x+20=0 to 2x^2-13x+20=0? How do you break the equation apart?

OpenStudy (anonymous):

sorry i meant to 2x^2-13x+20=0 to (x - 4)(2x - 5) = 0

OpenStudy (anonymous):

Broke it apart with common factors.

OpenStudy (anonymous):

x for 2x^2. 4 for 20.

OpenStudy (anonymous):

but 2,13 and 20 dont have anything in common?

OpenStudy (anonymous):

You do them by them self.

OpenStudy (anonymous):

Their own factor.

OpenStudy (anonymous):

so you factor 2x^2 by itself?

OpenStudy (anonymous):

In a way yes. By x.

OpenStudy (anonymous):

You want to factor it by x, because if you did whole number; you wouldn't get the correct answer.

OpenStudy (anonymous):

so you factor x out? im confused

OpenStudy (anonymous):

what type is it?

OpenStudy (anonymous):

as you see 2x^2 has two x's correct?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

x+x= 2x?

OpenStudy (anonymous):

You see the 2x is squared, correct?